OpenStudy (essancesosa_):

Write a quadratic equation with the given roots. Write the equation in the form , where a, b, and c are integers. -7 and -3

OpenStudy (rcp031):

We know that, standard form of quadratic equation is ax^2+bx+c Here, x=-7,x=-3 that means x+7=0 & x+3=0 (x+7)*(x+3)=0 x^2+7x+3x+21=0 x^2+10x+21=0 our solution.

OpenStudy (essancesosa_):

Thank you!